Mr Mangena invested an amount of R13890 divided in two different schemes A and B, at the simple interest rate of 14% per annum and 11% per annum respectively. If the total amount of simple interest earned in three years is R5 508. What was the amount invested in scheme B


1
Expert's answer
2021-05-17T16:57:52-0400

given information

invested amount = R13890

simple interest rate = 14%

scheme B simple interest = 11%

simple interest earned in three years = R5508


the invested amount in scheme A be RS x and that in scheme B be (R13890- x)

To compute the Sum invested in scheme B the expression will be

"= [X \times 14 \times3] 100 + [(R13890 -x) \times 11\times3] 100 = R5508"

"42x - 33x = R550800 -(R13890 \times 33)"

 9x = R550800 - R458370

  9x = R92430

"x = \frac{R92430 }{9}"

   x = R10270

therefore, the invested amount in scheme B is

   = R13890 - R10270

    = R3620